OpenStudy (hopehope118):

how would you solve: 4x/5 + 6x/10 = -12

OpenStudy (anonymous):

first of all, you would want to simplify 6x/10. and that becomes 3x/5 all write the rest of it down

OpenStudy (anonymous):

I'll *

hero (hero):

Yep, that's what I would do first as well. Smart approach to reduce the fraction first.

OpenStudy (anonymous):

\[\frac{ 4x }{ 5 }+\frac{ 3x }{ 5 }=-12\] multiply everything by 5 \[4x+3x=-60\] \[7x = -60\] \[x = \frac{ -60 }{ 7 }\]

hero (hero):

It's an improper fraction so reduce it

OpenStudy (hopehope118):

ok, thank you very much! :)

OpenStudy (anonymous):

sure